Solve the system using the substitution method. If the system has no solution or an infinite number of solutions, state this.
x+y=7
6x+6y=42

Found 3 solutions by Timnewman, ikleyn, MathTherapy:
Answer by Timnewman(323)

Solve the system using the substitution method. If the system has no solution or an infinite number of solutions, state this.
x+y=7
6x+6y=42
solution
x+y=7----(1)
6x+6y=42--(2)
divide equ 2 by 6
x+y=7---(3)
equate equ 1 and 3
x+y=x+y
x-y+x-y=0
0=0
from the above,the system has no solution.
TIMNEWMAN

Answer by ikleyn(52794)

.
Solve the system using the substitution method. If the system has no solution or an infinite number of solutions, state this.
x+y=7
6x+6y=42
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Divide the second equation by 6 (both sides),

and you will see that this equation become 

x + y = 7,

exactly as the first equation is.

So, you have here the system of two equivalent (depending) equations.


The system has infinitely many solutions.


Answer by MathTherapy(10552)


Solve the system using the substitution method. If the system has no solution or an infinite number of solutions, state this.
x+y=7
6x+6y=42
x + y = 7 ----- eq (i)
6x + 6y = 42 ------ eq (ii)
Multiply eq (i) by 6 to get the exact equation as eq (ii)
Divide eq (ii) by 6 to get the exact equation as eq (i)
This means the equations are the SAME, and therefore there is an .
However, you have to determine the answer using the SUBSTITUTION method. we then get:
x + y = 7____x = 7 - y ------ eq (i)
6x + 6y = 42 ------- eq (ii)
6(7 - y) + 6y = 42 -------- Substituting 7 - y for x in eq (ii)
42 - 6y + 6y = 42
- 6y + 6y = 42 - 42
0 = 0 ----- This is a TRUE statement and so, there is an: .